kekeman:

Determine how many, what type, and find the roots for f(x) = x4 + 21x2 − 100.

Rylee88:

x4 Find the variable (x) _4 + 21 = ____ * 2 = ____ - 100 = ___

kekeman:

Ok i was thinking that the roots were -x+2=0, x=-2, x-2=0, x=2, x-5i=0, x=5i, x+5i=0, x=-5

surjithayer:

it has four roots (being of 4th degree). put \[x^2=t,x^4=t^2\] f(x)=0 , gives \[t^2+21t-100=0\] \[t^2+25t-4t-100=0\] \[t(t+25)-4(t+25)=0\] \[(t+25)(t-4)=0\] \[(x^2+25)(x^2-4)=0\] \[(x^2-(-25))(x+2)(x-2)=0\] \[(x^2-25 \iota^2)(x+2)(x-2)=0\] \[(x+5 \iota)(x-5 \iota)(x+2)(x-2)=0\] \[x=5 \iota,-5 \iota,2,-2\]

kekeman:

@rylee88 wrote:
x4 Find the variable (x) _4 + 21 = ____ * 2 = ____ - 100 = ___
Ok i was thinking that the roots were -x+2=0, x=-2, x-2=0, x=2, x-5i=0, x=5i, x+5i=0, x=-5

kekeman:

@surjithayer wrote:
it has four roots (being of 4th degree). put \[x^2=t,x^4=t^2\] f(x)=0 , gives \[t^2+21t-100=0\] \[t^2+25t-4t-100=0\] \[t(t+25)-4(t+25)=0\] \[(t+25)(t-4)=0\] \[(x^2+25)(x^2-4)=0\] \[(x^2-(-25))(x+2)(x-2)=0\] \[(x^2-25 \iota^2)(x+2)(x-2)=0\] \[(x+5 \iota)(x-5 \iota)(x+2)(x-2)=0\] \[x=5 \iota,-5 \iota,2,-2\]
Thank you!

kekeman:

That was super helpful

surjithayer:

yw